OpenStudy (anonymous):

how to find all solutions of cot(3x) = − sqrt3 / 3 in interval [0,2pi]

OpenStudy (jdoe0001):

\(\bf cot(3x)=-\cfrac{\sqrt{3}}{3}\qquad take \quad cot^{-1} \\ \quad \\ cot^{-1}[cot(3x)]=cot^{-1}\left[-\cfrac{\sqrt{3}}{3}\right]\implies 3x=cot^{-1}\left[-\cfrac{\sqrt{3}}{3}\right]\)